04 July 2019: Lightsail 2 Mission Manager's Log

Article 5 Image

Spacecraft telemetry indicated that LightSail 2 continues to be in a good state of health. The flight team attempted to turn on gyros that are mounted on the payload interface board (PIB). The PIB gyros are higher fidelity than the gyros that are included as part of the Intrepid flight processor board. The PIB gyros did not respond to power-on commands; while the commands appeared to be received by the spacecraft, there was no increase in the power load and PIB gyros continued to show zero angular rates. Six test images were acquired from each of the two cameras mounted on the solar panels. The attitude control mode was changed from Mode 0 (detumble) to Mode 1 (Z-axis alignment). Mode 1 aligns the longitudinal (Z) axis of the CubeSat with the Earth’s magnetic field vector, which is a favorable attitude for communications. The command was successful, and the torque rod currents and Intrepid gyro angular rates changed in response. Notably, the Z-axis torque rod transitioned from a persistent zero reading to non-zero values, resolving a concern.

Troubleshooting of communication protocols fixed an ongoing problem with file downlinks, and the flight team was successful in downlinking stored telemetry files and thumbnail test images. A test image thumbnail is shown above. Overnight, the Cal Poly student operators continued to downlink stored telemetry files and began the downlink of a full resolution test image.

A LightSail 2 two-line element has not yet been provided by the 18th Space Control Squadron. The flight team continues to use the TLEs for object A.

Goal for today’s July 4 shift include a retry of the PIB gyro on command, and an autonomous momentum wheel test.

Anomaly List:
1. +Y solar panel shown as deployed in telemetry. Resolved: Beacons on July 3 showed all solar panels as stowed.
2. Spacecraft not responding as expected to file downlink commands. Resolved: Communications protocols update allowed multiple successful large file downlinks.
3. Z-axis torque rod showing zero current. Resolved: Transition to ADCS Mode 1 (Z-axis alignment) resulted in non-zero Z-axis torquer current.
4. Cal Poly (Marconi) tracking station unable to decode beacons, despite strong CW signals. Resolved: Ground system connections checked, and several beacons received in subsequent pass.
5. Solar panels indicate spacecraft in eclipse, ADCS algorithm flag (based on TLE) indicate illuminated. Open.